OpenStudy (anonymous):

Factor completely: 4x^2-8xy-3x+6y

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \sf 4x^2-8xy-3x+6y}\) \(\huge\color{blue}{ \sf 4x(x-2y)-3(x-2y)}\) can you do the last step?

OpenStudy (anonymous):

(4x-3)(x-2y) 4x-3=0 4x=3 x-3/4 x-2y=0 this is the one I'm stuck on????

OpenStudy (solomonzelman):

you are solving \(\color{blue}{ \sf 4x^2-8xy-3x+6y=0 }\) \(\color{blue}{ \sf 1 }\) equation \(\color{blue}{ \sf 2}\) unknown variables ?

OpenStudy (anonymous):

I have to have 2 answers and the x-2y=0 is how I'm suppose to get my second answer, but i don't know how to solve it.

OpenStudy (solomonzelman):

you can't solve 1 equation with 2 unknown variables !

OpenStudy (anonymous):

oh ok thank you!

OpenStudy (solomonzelman):

You welcome !